# Personalized Spiking Neural Networks with Ferroelectric Synapses for EEG Signal Processing

### v2

@@ 2.6 Device modeling and weight-update formulation
Fig. 3: Model a The change in normalized weight ($\Delta W$) is plotted with respect to the initial weight for long-term potentiation (LTP) and depression (LTD) including the device characterization data and the fitted model. b The evolution of weight with programming pulse number is plotted for the device data and model.
@@ 2.7 On-device learning
During online learning, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ using backpropagated gradients (Fig. 4). The hardware update is not performed as long as $|\delta w_{\mathrm{acc}}|<\varepsilon$, where $\varepsilon$ is a threshold that represents the minimum change that can be reliably programmed in the device. This accumulation-based update mechanism is conceptually similar to the mixed-precision learning strategy previously proposed for non-spiking networks with memristive [nandakumar2020mixed] and ferroelectric synapses [garg2025energy].
Once $|\delta w_{\mathrm{acc}}|\geq\varepsilon$, a single programming event is triggered. The sign of $\delta w_{\mathrm{acc}}$ determines whether a potentiation (LTP) or depression (LTD) pulse is applied to the active device. The resulting conductance change is computed from the Beta-kernel based device model. After the update, $\delta w_{\mathrm{acc}}$ is reset to zero. The updated device conductance $W$ is then mapped back to the neural weight domain through the scaling described earlier.
During the on-device learning phase, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ based on the backpropagated gradient. After each batch (64 trials), this accumulator is compared against asymmetric LTP/LTD thresholds, and the device update is chosen according to
@@
Here $\varepsilon$ is the base threshold, defined as a percentage of the fan-in-dependent weight range of each layer, and $\varepsilon_{\mathrm{asym}}$ scales the LTD threshold to reflect the larger step sizes observed for depression compared to potentiation in the device measurements.
Once an LTP or LTD programming event is triggered, the corresponding conductance change is computed from the beta-shaped kernel device model in Eq. (14). After the update, $\delta w_{\mathrm{acc}}$ is reset to zero, and the updated device conductance $W$ is mapped back to the neural weight domain via the scaling described earlier.
@@ 2.8 Subject specific transfer learning
Subjects excluded from each pretraining fold were the target subjects for fine tuning of the network. The pre-trained model, trained with the fully on-device learning method was used for subject-specific fine-tuning experiments. For each participant, the trials were divided into four equal subsets. In each fold, three subsets were used for fine-tuning, while the remaining subset was used solely for evaluation. Due to the fewer number of trials for each participant, and to emulate the online learning scenario, a batch size of one was used for this fine-tuning. Moreover, this fine-tuning was performed with the fitted memristor model. After cycling through all four folds, the predicted labels from each test subset were concatenated, and the overall classification accuracy for that participant was computed from the complete set of predictions.
@@ 3.1 Baseline
We first reproduced the floating-point training baseline of Kumar et al. [kumar2022decoding] using the same EEG motor-imagery dataset and network architecture, implemented within our spiking neural network framework. The validation accuracy from all five folds is plotted across training epochs in Fig. 5. The reproduced model achieved a test accuracy of 80.39 $\pm$ 2.98% for the two-class motor imagery task (left vs. right hand), closely matching the performance of 80.65 $\pm$ 3.83% reported earlier in  [kumar2022decoding] for spiking neural network. For non-spiking neural networks, [dose2018end] reported an accuracy of 82.43% using 4-s EEG segments, while [wang2020accurate] achieved 80.38% using the first 2 s of each trial. In contrast, as described in the Methods section, our approach considers only the first 1 s of the EEG segment for classification. This design choice is motivated by the need to substantially reduce intent-to-command latency, which is critical for real-time BCI operation.
@@ 3.2 On-device learning
Fig. 6: On-device learning a Validation accuracy of the network as a function of training epochs for different update thresholds ($\epsilon$) of 2.5%, 5%, and 7.5%. b Total number of weight updates accumulated across all synapses during training for the corresponding thresholds. c Validation accuracy plotted with respect to the total number of weight updates for different update thresholds. d Validation accuracy for asymmetric update thresholds ($\epsilon_{\mathrm{asym}}$) of 0.5, 1, and 2, where $\epsilon_{\mathrm{asym}}=\epsilon_{+}/\epsilon_{-}$ denotes the ratio between potentiation and depression thresholds.
@@
We trained the network from an untrained initialization using the threshold-based update triggering scheme, where the weight updates were computed using the fitted device model. The results for three different values of the update threshold ($\epsilon$) are shown in Fig. 6(a). In this experiment, the asymmetry factor was fixed to $\varepsilon_{asym}=1$ to isolate the effect of the relative threshold $\varepsilon$. A higher threshold corresponds to a coarser effective weight granularity, leading to less frequent device updates and slower convergence during training. Increasing the update threshold reduces the cumulative number of synaptic programming events required to reach a given classification accuracy, highlighting a trade-off between learning accuracy, convergence speed, and programming activity. As shown in Fig. 6(b) and (c), higher thresholds reach similar accuracy with fewer weight updates but require more learning iterations, resulting in lower programming energy and improved device lifetime. The test accuracies after training are $79.52\%\pm 3.15\%$, $79.13\%\pm 2.1\%$ and $78.91\%\pm 2\%$ for the threshold values of 2.5%, 5% and 7.5%, respectively.
@@
Fig. 6(d) further shows the impact of asymmetric update thresholds ($\epsilon_{\mathrm{asym}}$) on the validation accuracy. For this sweep, we set $\varepsilon=2.5\%$. The test accuracies are $79.92\%\pm 2.5\%$, $79.52\%\pm 3.15\%$, and $78.91\%\pm 2.76\%$ for asymmetry factors of 0.5, 1, and 2, respectively. $\epsilon_{\mathrm{asym}}$ of 0.5 shows the fastest convergence, and this is due to the fact that the LTD had steeper updates (A- = 0.033, A+ = 0.171), and thus this asymmetry in switching is compensated by the asymmetric thresholds. Nevertheless, the network exhibits convergence for all three degrees of asymmetry, showcasing the robustness of learning against asymmetric update behaviour. These results show that using a memristor model with a discrete number of levels and non-linear weight update, the accuracy of the network is comparable to the digital baseline.
@@ 3.3 Subject-specific transfer learning
EEG signals have a strong subject-specificity. In this section, we explore the feasibility of a subject-specific transfer learning (SSTL) strategy using ferroelectric synapses. In SSTL, a network pre-trained across four folds of participants and then subsequently fine-tuned for each individual of the fifth fold. The pre-trained model corresponds to a fully on-device trained model with $\epsilon$ of 2.5% and $\epsilon_{\mathrm{asym}}$ of 1, for a simpler implementation. These parameters and the memristive model was used for SSTL. Rather than retraining the full network, adaptation is restricted to selected subsets of layers in order to avoid underfitting due to a limited training data of only one subject. Fig. 7(a) compares the effect of retraining different layer groups using multiple learning rates. Transfer learning studies have shown that fine-tuning only higher-level layers often provides better generalization than retraining the full network, particularly when the target dataset is limited. Updating a large fraction of the network parameters can lead to overfitting or degradation of previously learned representations [yosinski2014transferable, li2017learning]. Based on this analysis, fine-tuning was restricted to the final fully connected layers (fc1 and fc2) using a learning rate of $6\times 10^{-4}$. The final test accuracy cumulated for all the participants, was 81.33%, corresponding to an improvement of 1.77% over the pre-trained network. Fig. 7 (b) shows the participant-wise test accuracy before and after subject-specific fine-tuning under this configuration. A systematic improvement in accuracy is observed for the majority of participants, despite a finite number of conductance levels for the ferroelectric weights and a non-linear update rule. This demonstrates that using ferroelectric synapses, personalization at the classifier level is sufficient to capture individual-specific EEG characteristics without modifying the full network.
@@ 3.4 Weight transfer and re-tuning
In this section, instead we investigate to which extent a transfer learning strategy can be implemented with ferroelectric synapses. The starting point is the baseline network (see Fig. 5) trained with floating-point precision. First, the effect of quantization on the network’s performance is measured. Synaptic weights were quantized to a specified number of discrete levels using uniform binning. For the two-class motor imagery classification task, as shown in Fig. 8 (a) the accuracy remains as high as 78% even when the weights are constrained to only three quantization levels ($\max$, 0, $\max$). This robustness to coarse weight quantization is a well-known strength [li2020robustness] of spiking neural networks, in which information is primarily encoded and propagated through spike timing rather than precise synaptic weight values [rueckauer2017conversion]. Consequently, the burden of precision is shifted from the weight domain to the temporal domain (e.g., spike times), enabling tolerance to low-resolution synaptic weights.
@@
To emulate device-level programming variability, Gaussian noise was added to the quantized weights. The noise standard deviation was defined as a fraction $\eta$ of the mean value of the non-zero quantized levels. The relationship between the network’s test accuracy and quantization levels was evaluated for $\eta$ ranging from 5% to 50%, the corresponding results are presented in Fig. 8(b). For the characterized device, the maximum $\eta$ of programmed weight from Fig. 2(e) is 3.75%. These measurements correspond to a single device and we can anticipate a higher programming error when accounting for device-to-device variability. Nevertheless, from the results, we can infer that only a drop in accuracy of 2% is observed for an extreme noise level of 25%. Prior work [shen2024conventional, stromatias2015robustness] has shown that SNNs can retain accuracy under substantial synaptic-weight noise and reduced precision, and in some regimes can be more robust to weight perturbations than comparable non-spiking neural networks [li2020robustness], supporting the use of low-resolution (e.g., low-bit or binary) weights for neuromorphic deployment.
@@
Fig. 9: Re-tuning of quantized weights. Validation accuracy as a function of training epochs for the pre-training phase, followed by quantization to three levels and additive noise with a standard deviation of 25% (a) and 50% (b), and subsequent re-tuning using the memristive device model over four epochs.
@@
Training a network entirely on hardware can significantly impact device endurance due to frequent programming operations. Therefore, we focus on a deployment strategy based on weight transfer followed by limited on-device re-tuning. In this approach, the network weights are first quantized and perturbed to account for programming variability, and the resulting network is subsequently fine-tuned for a small number of epochs using the memristive weight-update model. This setup emulates a practical open-loop deployment scenario, in which weights trained in software are transferred to a memristive crossbar and adapted on-device under realistic programming constraints. For three quantization levels and an additive noise level of $25\%$ and $50\%$, the validation accuracy recovers to approximately $80\%$ within a single training epoch, as shown in Fig. 9. The weight-update threshold in the accumulator was set to $\epsilon=2.5\%$. After re-tuning, we obtain a test accuracy of $79.08\%\pm 2.16\%$ for $\eta=25\%$ and $78.84\%\pm 2.39\%$ for $\eta=50\%$, which in both cases represents a slight decrease compared to the test accuracy of around $1.5\%$ after pre-training in floating-point precision.

### v3

@@ 2.6 Device modeling and weight-update formulation
Fig. 3: Model a The change in normalized weight ($\Delta W$) is plotted with respect to the initial weight for long-term potentiation (LTP) and depression (LTD) including the device characterization data and the fitted model. The dotted lines corresponds to the mean value of ($\Delta W$) obtained through the device characterization data for LTP and LTD. b The evolution of weight with programming pulse number is plotted for the device data and model.
@@ 2.7 On-device learning
During online learning, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ using backpropagated gradients (Fig. 4). The hardware update is not performed as long as $|\delta w_{\mathrm{acc}}|<\varepsilon$, where $\varepsilon$ is a threshold that represents the minimum change that can be reliably programmed in the device. This accumulation-based update mechanism is conceptually similar to the mixed-precision learning strategy previously proposed for non-spiking networks with memristive [42] and ferroelectric synapses [22].
Once $|\delta w_{\mathrm{acc}}|\geq\varepsilon$, a single programming event is triggered. The sign of $\delta w_{\mathrm{acc}}$ determines whether a potentiation (LTP) or depression (LTD) pulse is applied to the active device. The resulting conductance change is computed from the Beta-kernel based device model. After the update, $\delta w_{\mathrm{acc}}$ is reset to zero. The updated device conductance $W$ is then mapped back to the neural weight domain through the scaling described earlier. During the on-device learning phase, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ based on the backpropagated gradient. After each batch (64 trials), this accumulator is compared against asymmetric LTP/LTD thresholds, and the device update is chosen according to
@@
Here $\varepsilon$ is the base threshold, defined as a percentage of the fan-in-dependent weight range of each layer, and $\varepsilon_{\mathrm{asym}}$ scales the LTD threshold to reflect the larger step sizes observed for depression compared to potentiation in the device measurements. Once an LTP or LTD programming event is triggered, the corresponding conductance change is computed from the beta-shaped kernel device model in Eq. (14). After the update, $\delta w_{\mathrm{acc}}$ is reset to zero, and the updated device conductance $W$ is mapped back to the neural weight domain via the scaling described earlier.
@@ 2.8 Subject specific transfer learning
Subjects excluded from each pretraining fold were the target subjects for fine tuning of the network. The pre-trained model, trained with the fully on-device learning method was used for subject-specific fine-tuning experiments. For each participant, the trials were divided into four equal subsets. In each fold, three subsets were used for fine-tuning, while the remaining subset was used solely for evaluation. Due to the fewer number of trials for each participant, and to emulate the online learning scenario, a batch size of one was used for this fine-tuning. Moreover, this fine-tuning was performed with the fitted memristor model. After cycling through all four folds, classification accuracy was calculated for each fold, and the overall accuracy for the participant was obtained by averaging the accuracies across the four folds.
@@ 3.1 Baseline
We first reproduced the floating-point training baseline of Kumar et al. [32] using the same EEG motor-imagery dataset and network architecture, implemented within our spiking neural network framework. The validation accuracy from all five folds is plotted across training epochs in Fig. 5. The reproduced model achieved a test accuracy of 80.39 $\pm$ 2.98% for the two-class motor imagery task (left vs. right hand), closely matching the performance of 80.65 $\pm$ 3.83% reported earlier in  [32] for spiking neural network. For non-spiking neural networks, [13] reported an accuracy of 82.43% using 4-s EEG segments, while [62] achieved 80.38% using the first 2 s of each trial. In contrast, as described in the Methods section, our approach considers only the first 1 s of the EEG segment for classification. This design choice is motivated by the need to substantially reduce intent-to-command latency, which is critical for real-time BCI operation.
@@ 3.2 On-device learning
Fig. 6: On-device learning a Validation accuracy of the network as a function of training epochs for different update thresholds ($\epsilon$) of 2.5%, 5%, and 7.5%. b Total number of weight updates accumulated across all synapses during training for the corresponding thresholds. c Validation accuracy plotted with respect to the total number of weight updates for different update thresholds. d Validation accuracy for asymmetric update thresholds ($\epsilon_{\mathrm{asym}}$) of 0.5, 1, and 2, where $\epsilon_{\mathrm{asym}}=\epsilon_{+}/\epsilon_{-}$ denotes the ratio between potentiation and depression thresholds. e Validation accuracy during pre-training when cycle-to-cycle variability is included in the device-update model by sampling the update amplitude ($A_{+}$ and $A_{-}$) with a relative standard deviation of 5%, compared with the nominal case without variability.
@@
We trained the network from an untrained initialization using the threshold-based update triggering scheme, where the weight updates were computed using the fitted device model. The results for three different values of the update threshold ($\epsilon$) are shown in Fig. 6(a). In this experiment, the asymmetry factor was fixed to $\varepsilon_{asym}=1$ to isolate the effect of the relative threshold $\varepsilon$. The threshold values of 2.5%, 5%, and 7.5% were chosen to lie in the same range as the experimentally observed programming granularity of the ferroelectric device. From the fitted device-update characteristics (see Fig. 3a), the programming increments span from below 1% up to approximately 10% for LTP and 15% for LTD depending on the conductance state, while the mean update magnitudes are around 4% for LTP and 5% for LTD. The selected thresholds therefore probe a practically relevant regime in which the triggering condition is comparable to the typical device programming step size. A higher threshold corresponds to a coarser effective weight granularity, leading to less frequent device updates and slower convergence during training. Increasing the update threshold reduces the cumulative number of synaptic programming events required to reach a given classification accuracy, highlighting a trade-off between learning accuracy, convergence speed, and programming activity. As shown in Fig. 6(b) and (c), higher thresholds reach similar accuracy with fewer weight updates but require more learning iterations, resulting in lower programming energy and improved device lifetime. The test accuracies after training are $79.52\%\pm 3.15\%$, $79.13\%\pm 2.1\%$ and $78.91\%\pm 2\%$ for the threshold values of 2.5%, 5% and 7.5%, respectively.
@@
Fig. 6(d) further shows the impact of asymmetric update thresholds ($\epsilon_{\mathrm{asym}}$) on the validation accuracy. For this sweep, we set $\varepsilon=2.5\%$. The test accuracies are $79.92\%\pm 2.5\%$, $79.52\%\pm 3.15\%$, and $78.91\%\pm 2.76\%$ for asymmetry factors of 0.5, 1, and 2, respectively. $\epsilon_{\mathrm{asym}}$ of 0.5 shows the fastest convergence, and this is due to the fact that the LTD had steeper updates (A- = 0.033, A+ = 0.171), and thus this asymmetry in switching is compensated by the asymmetric thresholds. Nevertheless, the network exhibits convergence for all three degrees of asymmetry, showcasing the robustness of learning against asymmetric update behaviour. These results show that using a memristor model with a discrete number of levels and non-linear weight update, the accuracy of the network is comparable to the digital baseline. Furthermore, to account for cycle-to-cycle variability representative of the physical device, we sampled the update amplitude in the Beta-shaped kernel from a normal distribution with a relative standard dispersion of 5%. From Fig. 2(f), the programmed weight variation in the characterized device reaches a maximum relative standard dispersion of 3.75%. In this way, each programming event during training experiences a slightly different effective update magnitude, thereby emulating cycle-to-cycle fluctuations in the ferroelectric synapse. As shown in Fig. 6(e), the convergence trajectory remains very similar to the nominal case, and the final test accuracy after 20 epochs of training reaches $79.60\%\pm 2.92\%$ with cycle-to-cycle variations included. This indicates the robustness of the proposed training framework to stochastic programming variations observed in the device.
@@ 3.3 Subject-specific transfer learning
EEG signals have a strong subject-specificity. In this section, we explore the feasibility of a subject-specific transfer learning (SSTL) strategy using ferroelectric synapses. In SSTL, a network pre-trained across four folds of participants and then subsequently fine-tuned for each individual of the fifth fold. The pre-trained model corresponds to a fully on-device trained model with $\epsilon$ of 2.5% and $\epsilon_{\mathrm{asym}}$ of 1, for a simpler implementation. These parameters and the memristive model was used for SSTL. Rather than retraining the full network, adaptation is restricted to selected subsets of layers in order to avoid underfitting due to a limited training data of only one subject. Fig. 7(a) compares the effect of retraining different layer groups using multiple learning rates. Transfer learning studies have shown that fine-tuning only higher-level layers often provides better generalization than retraining the full network, particularly when the target dataset is limited. Updating a large fraction of the network parameters can lead to overfitting or degradation of previously learned representations [64, 35]. Based on this analysis, fine-tuning was restricted to the final fully connected layers (fc1 and fc2) using a learning rate of $6\times 10^{-4}$. The final test accuracy cumulated for all the participants, was 81.33%, corresponding to an improvement of 1.77% over the pre-trained network. Fig. 7 (b) shows the participant-wise test accuracy before and after subject-specific fine-tuning under this configuration. A systematic improvement in accuracy is observed for the majority of participants, despite a finite number of conductance levels for the ferroelectric weights and a non-linear update rule. This demonstrates that using ferroelectric synapses, personalization at the classifier level is sufficient to capture individual-specific EEG characteristics without modifying the full network.
@@ 3.4 Weight transfer and re-tuning
In this section, instead we investigate to which extent a transfer learning strategy can be implemented with ferroelectric synapses. The starting point is the baseline network (see Fig. 5) trained with floating-point precision. First, the effect of quantization on the network’s performance is measured. Synaptic weights were quantized to a specified number of discrete levels using uniform binning. For the two-class motor imagery classification task, as shown in Fig. 8 (a) the accuracy remains as high as 78% even when the weights are constrained to only three quantization levels ($\max$, 0, $\max$). This robustness to coarse weight quantization is a well-known strength [34] of spiking neural networks, in which information is primarily encoded and propagated through spike timing rather than precise synaptic weight values [48]. Consequently, the burden of precision is shifted from the weight domain to the temporal domain (e.g., spike times), enabling tolerance to low-resolution synaptic weights.
@@
To emulate device-level programming variability, Gaussian noise was added to the quantized weights. The noise standard deviation was defined as a fraction $\eta$ of the mean value of the non-zero quantized levels. The relationship between the network’s test accuracy and quantization levels was evaluated for $\eta$ ranging from 5% to 50%, the corresponding results are presented in Fig. 8(b). For the characterized device, the maximum $\eta$ of programmed weight from Fig. 2(f) is 3.75%. These measurements correspond to a single device and we can anticipate a higher programming error when accounting for device-to-device variability. Nevertheless, from the results, we can infer that only a drop in accuracy of 2% is observed for an extreme noise level of 25%. Prior work [51, 56] has shown that SNNs can retain accuracy under substantial synaptic-weight noise and reduced precision, and in some regimes can be more robust to weight perturbations than comparable non-spiking neural networks [34], supporting the use of low-resolution (e.g., low-bit or binary) weights for neuromorphic deployment. While these results highlight the robustness of the adopted SNN to coarse weight quantization and programming noise, a direct comparison with a non-spiking network under the same architectural and ferroelectric hardware-mapping constraints remains an important direction for future work.
@@
Fig. 9: Re-tuning of quantized weights. Validation accuracy across epochs is shown for the pre-training phase, followed by quantization to three levels and additive noise with standard deviation $\eta=25\%$ (a) and $\eta=50\%$ (d), and subsequent re-tuning using the memristive device model. Panels (b) and (e) show the re-tuning phase extended to 10 epochs, while panels (c) and (f) report the cumulative number of weight updates during re-tuning for the corresponding noise levels.
@@
Training a network entirely on hardware can significantly impact device endurance due to frequent programming operations. Therefore, we focus on a deployment strategy based on weight transfer followed by limited on-device re-tuning. In this approach, the network weights are first quantized and perturbed to account for programming variability, and the resulting network is subsequently fine-tuned for a small number of epochs using the memristive weight-update model. This setup emulates a practical open-loop deployment scenario, in which weights trained in software are transferred to a memristive crossbar and adapted on-device under realistic programming constraints. For three quantization levels and an additive noise level of $25\%$ and $50\%$, the validation accuracy recovers to approximately $80\%$ within a single training epoch, as shown in Fig. 9. The weight-update threshold in the accumulator was set to $\epsilon=2.5\%$. After re-tuning, we obtain a test accuracy of $79.08\%\pm 2.16\%$ for $\eta=25\%$ and $78.84\%\pm 2.39\%$ for $\eta=50\%$, which in both cases represents only a slight decrease of around 1.5 percentage points relative to the test accuracy after pre-training in floating-point precision. An extended re-tuning the network for 10 epochs yields only limited additional improvement compared with the results obtained after 4 epochs. For $\eta=25\%$, the test accuracy increases only slightly from $79.08\%\pm 2.16\%$ to $80.46\%\pm 2.63\%$, and for $\eta=50\%$, from $78.84\%\pm 2.39\%$ to $79.50\%\pm 2.91\%$. Rather than the rapid recovery observed during the first one to two re-tuning epochs, the validation accuracy fluctuates around a similar level after the fourth epoch. Meanwhile, the cumulative number of weight updates still increases, but progressively more slowly, indicating diminishing returns in accuracy relative to additional programming activity. This supports the choice of limiting the re-tuning phase to only a few epochs.
3.5 Hardware implications
Beyond the algorithmic results, the adopted bipolar weight mapping also has important hardware implications. In the present scheme, signed weights are represented using a differential-pair configuration in which one device is fixed at a mid-level conductance and the other encodes the effective weight. This differs from conventional differential encoding, commonly used in phase-change memory (PCM) [43] and other device technologies, where one device stores the positive and the other the negative contribution of a weight, often motivated by strongly asymmetric potentiation and depression characteristics. In contrast, ferroelectric devices exhibit granular and bidirectional conductance modulation, enabling both LTP and LTD on a single device. This alleviates conductance saturation and reduces the refresh operations often required in two-device differential schemes [42]. In addition, this approach requires only one column of devices as a reference column rather than duplicating the full array. This reduces hardware overhead compared with fully differential signed-weight mappings, in which two programmable devices or columns are required per weight. In the present scheme, the remaining cost is mainly associated with the reference column and the corresponding subtraction or readout circuitry, rather than a second fully programmable branch. As a result, both area and programming energy can be reduced relative to conventional two-device differential implementations while still supporting bipolar weight representation. Offset-based bipolar synapses have also been demonstrated in hardware using digital offset correction [39], in online learning with magnetic tunnel junctions [41], and through reference columns to mitigate sneak-path effects [16].
[…]
In contrast to the synaptic model, which is tied to experimentally characterized device behavior, the neuron in this study is treated at the level of functional modeling rather than circuit implementation. The discrete-time LIF dynamics are therefore used as an abstract computational model and do not imply a specific circuit-level realization. Rather, the LIF equations define the temporal behavior of the network, whereas the hardware contribution of this study focuses on the ferroelectric synaptic devices used to store and update the network weights under realistic programming constraints. From a deployment perspective, this formulation is compatible with mixed-signal implementations [21] in which ferroelectric devices implement programmable synaptic memory, while neuron-state updates such as integration, leakage, reset, and recurrence are handled either digitally or via custom analog neuron circuits.
In this work, we adopt surrogate-gradient descent with threshold-triggered programming as a device-aware training interface, rather than as a claim of fully hardware-native synaptic learning. This choice is motivated by the target application: EEG motor-imagery decoding is a supervised classification problem, for which gradient-based optimization provides a direct and well-established way to optimize task-level accuracy in deep spiking networks. Our goal here is therefore not to propose a new local ferroelectric learning rule, but to evaluate how supervised SNN training and post-deployment personalization can be constrained by experimentally measured ferroelectric programming dynamics.
The thresholded accumulation scheme should therefore be interpreted as an interface between supervised gradient-based optimization and the physical programming constraints of the ferroelectric synapse, rather than as a replacement of the device dynamics by a purely digital learning rule. The threshold determines only when a programming event is invoked; the actual conductance update is still obtained from the fitted beta-kernel device model and thus reflects the experimentally measured state- and polarity-dependent response of the device.
3.6 Energy and feasibility
To provide a first-order indication of the potential energy efficiency of the proposed approach, we estimated the inference and programming energy using measured device characteristics together with literature-reported values for ultra-low-power neuron circuits.
During inference, the mean effective MAC count per sample is 1,466,355 when averaged over 45 trials for subject 1, due to sparse activity. Using the measured maximum conductance of the characterized device, $G_{\max}=6.5\times 10^{-8}~\mathrm{S}$, together with a read voltage of $100~\mathrm{mV}$ and an integration time of $500~\mu\mathrm{s}$, the worst case synaptic energy per MAC can be estimated as
$E_{\mathrm{MAC}}\approx V_{\mathrm{read}}^{2}Gt_{\mathrm{int}}=(0.1)^{2}\times 6.5\times 10^{-8}\times 5\times 10^{-4}=3.25\times 10^{-13}~\mathrm{J}=0.325~\mathrm{pJ}.$ (16)
This yields an estimated synaptic inference energy of
$E_{\mathrm{syn}}=1{,}466{,}355\times 0.325~\mathrm{pJ}\approx 4.77\times 10^{-7}~\mathrm{J}=0.477~\mu\mathrm{J}.$ (17)
We further estimated the neuron energy using the ultra-low-power neuron reported in [55], assuming $4~\mathrm{fJ}$ per spike dynamic energy and $100~\mathrm{pW}$ static power:
$E_{\mathrm{neuron}}=P_{\mathrm{static}}\,T_{\mathrm{sample}}\,N_{\mathrm{neurons}}+E_{\mathrm{spike}}\,N_{\mathrm{spikes}}.$ (18)
Using a mean spike count of 569,438 over a $1~\mathrm{s}$ sample and a total of 11,010 neurons, the static neuron energy is estimated to be $1.101~\mu\mathrm{J}$, and the dynamic neuron energy is estimated to be $0.0023~\mu\mathrm{J}$, resulting in a total neuron energy of $1.103~\mu\mathrm{J}$. The total estimated inference energy per sample is therefore
$E_{\mathrm{inf}}=E_{\mathrm{syn}}+E_{\mathrm{neuron}}\approx 1.580~\mu\mathrm{J},$ (19)
and is dominated by the static neuron contribution. As a reference point, Kumar et al. reported an inference energy of $3.60~\mathrm{mJ}$ on Intel’s Loihi for a reduced version of the same broad SNN architecture, without adaptive thresholds or trainable decay parameters due to hardware constraints [33]. While this value provides a useful system-level hardware reference, it is not directly comparable to the present estimate, which is derived from literature-reported device- and circuit-level figures of merit rather than from a fabricated end-to-end implementation, and does not include peripheral, communication, memory-access, or control overheads. While this comparison is only indicative, it is consistent with the broader promise of analog in-memory computing for energy-efficient neuromorphic processing.
Finally, for device programming, we estimate from Fig. 6(c) that the total number of weight-update events is on the order of $10^{5}$ for an update threshold of $7.5\%$. From the JJVV sweeps in Fig. 2(c), the maximum absolute current density is $0.2660~\mathrm{A/cm^{2}}$ at a programming voltage of $1.25~\mathrm{V}$. Using the device area of $8300~\mu\mathrm{m}^{2}$ and a programming pulse width of $500~\mu\mathrm{s}$, the programming energy can be estimated as
$E_{\mathrm{pulse}}=JAVt_{p}$ (20)
which yields an energy per programming pulse of approximately $13.8~\mathrm{nJ}$, corresponding to a total programming energy of approximately $1.38\times 10^{-3}~\mathrm{J}$ ($1.38~\mathrm{mJ}$) for $10^{5}$ update events. However, this estimate corresponds to the present measurement conditions, namely a device area of $8300~\mu\mathrm{m}^{2}$ and a programming pulse width of $500~\mu\mathrm{s}$, and is therefore expected to be substantially higher than in scaled devices operated with shorter pulses. Recent work has shown that, by scaling ferroelectric hafnia devices below $100~\mu\mathrm{m}^{2}$, reliable programming with $20$ ns pulses can be achieved with a maximum energy of about $3$ pJ per update [3]. These results indicate that the programming energy of the present device could be considerably reduced by lateral scaling and shorter pulse operation.
These estimates account only for the electrical energy dissipated in the device during inference and programming and do not include the energy required for gradient computation, peripheral circuitry, or the generation of programming pulses, for example, through a DAC. Moreover, the adaptation considered here is not intended as a lifelong continuous-learning scenario, but rather as a one-time or infrequent post-deployment personalization step, so the associated programming cost should be interpreted as a bounded adaptation overhead. Overall, we view the present study as a validation at the algorithmic and device-aware modeling level, while a full end-to-end hardware implementation and direct system-level measurement of inference and adaptation energy, including peripheral and control overheads, remain important directions for future work.
3.7 Outlook
A complete hardware-aware deployment study should also account for crossbar-level non-idealities beyond the device-update characteristics considered here, including IR drop, sneak-path currents, and half-select disturb. These effects were not explicitly modeled in the present work, but in practical implementations they can be mitigated through bounded array dimensions, selector-based cells such as 1T1R or 1S1R, and appropriate write/verify schemes [29]. In addition, such effects can be incorporated at the algorithm level through hardware-aware training, calibration, and post-programming fine-tuning using realistic crossbar models [46]. Extending the present framework in this direction is an important next step toward fully hardware-issues-aware neuromorphic processing.
An important next step is to move from supervised calibration toward lifelong adaptation that does not depend on continuous labeled supervision. In this perspective, a practical deployment pathway is to use labeled data to obtain an accurate initialization, as done in this work, followed by low-overhead online adaptation through local, hardware-native learning rules. In particular, three-factor learning with eligibility traces [23, 17] offers a hardware-compatible alternative to backpropagation through time by combining local synaptic traces with a global modulatory signal. Ferroelectric synapses have also demonstrated plasticity behaviors that can support Hebbian learning [20], suggesting a feasible route toward device-native learning in future systems.